OpenStudy (howard-wolowitz):

Urgent help only 11 question and I'll medal!

OpenStudy (howard-wolowitz):

Does anyone know?

OpenStudy (howard-wolowitz):

@mathmate

OpenStudy (michele_laino):

\[f\left( 0 \right) = \frac{{4 \cdot 0 + 12}}{{0 - 4}} = ...?\]

OpenStudy (michele_laino):

sorry is it your homework?

OpenStudy (howard-wolowitz):

\[4x ^{2}\]

OpenStudy (michele_laino):

it is not among your options, please

OpenStudy (howard-wolowitz):

I already have D picked

OpenStudy (michele_laino):

I'm sorry but option D is wrong

OpenStudy (howard-wolowitz):

ohh ok so it does have \[4x ^{2}+12\]

OpenStudy (michele_laino):

yes!

OpenStudy (howard-wolowitz):

So I think that i messed up with my denominator then right?

OpenStudy (michele_laino):

I think that the denominator is right

OpenStudy (howard-wolowitz):

oh so I just got my sign wrong then so it would be A then

OpenStudy (michele_laino):

that's right!